# Anatomy of critical fluctuations in hadronic matter

### v1

@@ II Parity doublet model
In this work, we focus on a qualitative description of the fluctuations of the net-baryon number density. We consider a representative value of $m_{0}^{N}=0.8~{}\rm GeV$. This large value is motivated by the recent lattice QCD results [25, 26, 24]. Similar values have also been used in the literature [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 63, 46, 47, 48, 81, 49, 50, 51, 52, 53, 82, 54, 55, 56, 62]. Note, however, that the qualitative structure of the results presented in this paper does not depend on the choice of $m_{0}^{N}$. There are two remaining parameters, $m_{0}^{\Delta}$ and $r_{\Delta}$. At vanishing baryon chemical potential, the EOS does not depend on $r_{\Delta}$ due to the vanishing $\omega$ at $\mu_{B}=0$#1#1#1We note that while the EOS at vanishing baryon chemical potential and finite temperature, i.e., $p(\epsilon)$ is not affected by the values of repulsive couplings, the fluctuations depend on them. See, for example, [64].. Therefore, we first use the finite-temperature EOS at $\mu_{B}=0$ to determine $m_{0}^{\Delta}$. We find that for the chosen value of $m_{0}^{N}$, the model yields a smooth chiral crossover transition for $m_{0}^{\Delta}\gtrsim 0.95~{}\rm GeV$. We also require $m_{0}^{N}\lesssim m_{0}^{\Delta}$. In general, too low values of $m_{0}^{\Delta}$ lead to a first-order phase transition at $\mu_{B}=0$, and also lead to the appearance of $\Delta$ matter at subsaturation densities; thus, spoiling the ground state properties [65]. On the other hand, setting $m_{0}^{\Delta}\rightarrow\infty$ suppresses the $\Delta$ states, and the EOS effectively corresponds to the pure nucleonic EOS. In the present work, we choose $m_{0}^{\Delta}=0.95~{}\rm GeV$ because it yields the lowest temperature of the chiral crossover transition at the vanishing baryon chemical potential. In Fig. 1, we plot the normalized expectation value of the $\sigma$ mean field at vanishing baryon chemical potential for the pure nucleonic (no $\Delta$) and $m_{0}^{\Delta}=0.95~{}$GeV models. Including $\Delta$ lowers the pseudocritical temperature from $T_{c}\simeq 0.21~{}$GeV to $T_{c}\simeq 0.17~{}$GeV. With $m_{0}^{\Delta}$ fixed, we are left with a single parameter, $r_{\Delta}$, and we systematically study the influence of repulsion on the fluctuations of the net-baryon number density. The phase structure at vanishing baryon chemical potential is independent of the strength of the repulsive couplings. Therefore, the strongest dependence of the EOS on $r_{\Delta}$ is expected at low temperature and high density. In Fig. 2, we show the phase diagram of the parity doublet model for $r_{\Delta}=0.8,~{}1.2$. The liquid-gas crossover line obtained at a minimum of $\partial\sigma/\partial\mu_{N_{+}}$ (see Sec. III for details) is almost independent of the choice of $r_{\Delta}$. This is to be expected since the parameters are chosen so as not to affect the properties of the nuclear ground state at low temperatures and $r_{\Delta}$ does not affect the EOS at vanishing baryon chemical potential. Other lines, obtained from the corresponding inflection points of $\partial\sigma/\partial\mu_{\alpha}$, are sensitive to the value of $r_{\Delta}$ at low temperatures. They converge and continue almost as a single line with increasing temperature until $T_{c}$ is reached at vanishing baryon chemical potential. In this work, we pay particular attention to the low temperature and high baryon chemical potential part of the phase diagram. We systematically study the influence of $\Delta$ matter on the fluctuations of the net-baryon number in dense matter.
@@ III Susceptibilities of the net-baryon number density
which contain contributions from the nucleonic, $\Delta$ sectors, and terms that mix them. We note that in [64], a simplified system consisting of a single parity doublet was considered. By explicitly including the $\Delta_{\pm}$ states, besides the individual nucleonic and $\Delta$ sectors, four additional susceptibilities that mix them have to be considered. The individual terms in Eq. (22) equation are given as follows
@@
The scaled variances, $R_{2,1}^{\alpha\beta}$, can be defined also for cumulants related to the nucleonic and $\Delta$ sectors in Eq. (24). We note that, in general, the partial scaled variances $R_{2,1}^{\alpha\beta}$, are not additive. For example, $R_{2,1}^{NN}+R_{2,1}^{\Delta\Delta}+2R_{2,1}^{N\Delta}\neq R_{2,1}^{B}$. We note that due to experimental limitations, the net-proton number fluctuations are assumed to reflect the fluctuations of the net-baryon number. However, this relation has not been yet explored in theoretical models with dynamical chiral symmetry restoration.
@@ IV Results
First, we study the susceptibilities at finite temperatures and vanishing baryon chemical potential. They are shown in Fig. 4. It is instructive to compare our results with those of the truncated hadron resonance gas (tHRG) model, where the only degrees of freedom are nucleons, $\Delta$s, and their chiral partners. The HRG model is widely used for the description of matter under extreme conditions, e.g., in the context of heavy-ion collision phenomenology [85, 86, 87, 88, 89, 90]. The thermodynamic potential of the tHRG model is a mixture of uncorrelated ideal gases of stable particles:
@@
Now we turn to the finite baryon chemical potential regime and consider fixed $T=0.03~{}\rm GeV$. In Fig. 5, we show the net-baryon number susceptibility. The behavior at $\mu_{B}<1~{}\rm GeV$ is similar for all the values of $r_{\Delta}$ shown in the figure. The curves show a well-pronounced peak which is a remnant of the liquid-gas phase transition. Small deviations from the pure nucleonic result are due to a small fraction of the $\Delta$ matter present at $T=0.03~{}$GeV. At higher chemical potentials, the susceptibilities develop plateaux. A small peak is seen for the pure nucleonic EOS around $\mu_{B}=1.4~{}\rm GeV$, which is a remnant of the first-order chiral phase transition at low temperatures. The secondary peak for $r_{\Delta}=0.8$ is a result of lower repulsion for the $\Delta$ matter, leading to an early onset of the $\Delta_{+}$ state. The result for $r_{\Delta}=1.1$ is very close to the pure nucleonic EOS. We note that for $r_{\Delta}=1.2$, the net-baryon number susceptibility converges to that of pure nucleonic EOS. This convergence is due to the suppression of $\Delta$ matter by the repulsive interactions. However, the net-baryon number susceptibility at higher baryon chemical potential shows very little sensitivity to the choice of $r_{\Delta}$ and it depends only on the asymptotic matter composition, as discussed in the previous section.
The individual susceptibilities $\chi_{2}^{\alpha\beta}$ also depend on the values of repulsive couplings. This can be seen in the right panel of Fig. 6. For $r_{\Delta}<1$, the susceptibility is dominated by the contribution of $\Delta_{\pm}$. Similarly, for $r_{\Delta}>1$, they are dominated by the contribution of $N_{\pm}$. This is to be expected from the relative abundances shown in Fig. 3 and discussed earlier. For $r_{\Delta}=1$ all susceptibilities are not negligible. In general, the susceptibilities $\chi_{2}^{\alpha\alpha}$ at high baryon chemical potential are much larger than the corresponding peak from the liquid-gas transition below $\mu_{B}=1~{}$GeV. At the same time, the positive-negative parity correlations within the same doublet become negative and show minima of similar size as the peaks in $\chi_{2}^{\alpha\alpha}$. The correlations between $N_{\pm}$ and $\Delta_{\pm}$ are shown in the left panel of Fig. 6 and have a qualitatively similar structure for different values of $r_{\Delta}$. However, their magnitude depends on the choice of the repulsive coupling and they are non-negligible, except for $r_{\Delta}=1.2$, for which the $\Delta$ matter is almost completely suppressed. Interestingly, we find that the susceptibilities of the opposite parity states (which are not chiral partners), i.e., $\chi_{2}^{N_{+}\Delta_{-}}$ and $\chi_{2}^{N_{-}\Delta_{+}}$ become large and negative. They become dominant and have minima around the chiral crossover, providing an additional source of cancellation of the net-baryon number fluctuations at finite density.
@@
As shown in Fig. 8, qualitatively similar structures can be seen in the scaled variances defined in Eqs. (27). At low chemical potential, the sensitivity to the liquid-gas transition is seen in $R_{2,1}^{N_{+}N_{+}}$. At high baryon chemical potential as the chiral symmetry becomes restored, the signal is dominated by the scaled variances of the negative parity states, which develop peaks related to the onset of the chiral partners, $N_{-}$ and $\Delta_{-}$. The mixed scaled variances between $N_{\pm}$ and $\Delta_{\pm}$ show patterns similar to the susceptibilities $\chi_{2}^{\alpha\beta}$. At high baryon chemical potential, the ratios $R_{2,1}^{N_{+}\Delta_{+}}$ and $R_{2,1}^{N_{-}\Delta_{-}}$ feature peaks related to the restoration of chiral symmetry while $R_{2,1}^{N_{-}\Delta_{+}}$ and $R_{2,1}^{N_{+}\Delta_{-}}$ feature minima.
Most of the structure seen in the scaled variances is washed out when the full nucleonic and $\Delta$ sectors are considered as defined in Eqs. (22) and (27). This is shown in Fig. 9. For $r_{\Delta}=0.8$, $R_{2,1}^{NN}\simeq R_{2,1}^{B}$ below $\mu_{B}\simeq 0.9~{}\rm GeV$ until $\Delta$ matter becomes populated. This is also reflected in the enhancement of $R_{2,1}^{\Delta\Delta}$, which converges to $R_{2,1}^{B}$ with a higher baryon chemical potential. At the same time $R_{2,1}^{NN}$ goes to unity, due to $r_{\Delta}<1$. Similarly for $r_{\Delta}=1.0$, $R_{2,1}^{NN}$ deviates from $R_{2,1}^{B}$ as soon as the $\Delta$ matter sets in, i.e., $R_{2,1}^{\Delta\Delta}$ deviates from unity. However, since at high density the composition is determined by the degeneracy factors, none of the ratios converges to $R_{2,1}^{B}$. For $r_{\Delta}=1.2$, $R_{2,1}^{\Delta\Delta}\simeq 1$, due to the suppression of $\Delta_{\pm}$ states. Consequently, $R_{2,1}^{NN}\simeq R_{2,1}^{B}$. However, $R_{2,1}^{B}$ is suppressed compared to other scaled variances, regardless of the choice of $r_{\Delta}$.

### v2

@@ II Parity doublet model
In this work, we focus on a qualitative description of the fluctuations of the net-baryon number density. We consider a representative value of $m_{0}^{N}=0.8~{}\rm GeV$. This large value is motivated by the recent lattice QCD results [25, 26, 24]. Similar values have also been used in the literature [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 63, 46, 47, 48, 81, 49, 50, 51, 52, 53, 82, 54, 55, 56, 62]. Note, however, that the qualitative structure of the results presented in this paper does not depend on the choice of $m_{0}^{N}$. There are two remaining parameters, $m_{0}^{\Delta}$ and $r_{\Delta}$. At vanishing baryon chemical potential, the EOS does not depend on $r_{\Delta}$ due to the vanishing $\omega$ at $\mu_{B}=0$#1#1#1We note that while the EOS at vanishing baryon chemical potential and finite temperature, i.e., $p(\epsilon)$ is not affected by the values of repulsive couplings, the fluctuations depend on them. See, for example, [64].. Therefore, we first use the finite-temperature EOS at $\mu_{B}=0$ to determine $m_{0}^{\Delta}$. We find that for the chosen value of $m_{0}^{N}$, the model yields a smooth chiral crossover transition for $m_{0}^{\Delta}\gtrsim 0.95~{}\rm GeV$. We also require $m_{0}^{N}\lesssim m_{0}^{\Delta}$. In general, too low values of $m_{0}^{\Delta}$ lead to a first-order phase transition at $\mu_{B}=0$, and also lead to the appearance of $\Delta$ matter at subsaturation densities; thus, spoiling the ground state properties [65]. On the other hand, setting $m_{0}^{\Delta}\rightarrow\infty$ suppresses the $\Delta$ states, and the EOS effectively corresponds to the pure nucleonic EOS. In the present work, we choose $m_{0}^{\Delta}=0.95~{}\rm GeV$ because it yields the lowest temperature of the chiral crossover transition at the vanishing baryon chemical potential. In Fig. 1, we plot the normalized expectation value of the $\sigma$ mean field at vanishing baryon chemical potential for the pure nucleonic (no $\Delta$) and $m_{0}^{\Delta}=0.95~{}$GeV models. Including $\Delta$ lowers the pseudocritical temperature from $T_{c}\simeq 0.21~{}$GeV to $T_{c}\simeq 0.17~{}$GeV. With $m_{0}^{\Delta}$ fixed, we are left with a single parameter, $r_{\Delta}$, and we systematically study the influence of repulsion on the fluctuations of the net-baryon number density. The phase structure at vanishing baryon chemical potential is independent of the strength of the repulsive couplings. Therefore, the strongest dependence of the EOS on $r_{\Delta}$ is expected at low temperature and high density.
In Fig. 2, we show the phase diagram of the parity doublet model for $r_{\Delta}=0.8,~{}1.2$. The liquid-gas crossover line obtained at a minimum of $\partial\sigma/\partial\mu_{N_{+}}$ (see Sec. III for details) is almost independent of the choice of $r_{\Delta}$. This is to be expected since the parameters are chosen so as not to affect the properties of the nuclear ground state at low temperatures and $r_{\Delta}$ does not affect the EOS at vanishing baryon chemical potential. Other lines, obtained from the corresponding inflection points of $\partial\sigma/\partial\mu_{\alpha}$, are sensitive to the value of $r_{\Delta}$ at low temperatures. They converge and continue almost as a single line with increasing temperature until $T_{c}$ is reached at vanishing baryon chemical potential. In this work, we pay particular attention to the low temperature and high baryon chemical potential part of the phase diagram. We systematically study the influence of $\Delta$ matter on the fluctuations of the net-baryon number in dense matter. We remark that the liquid-gas crossover transition develops a critical point at $T\simeq 0.016~{}$GeV and turns into an ordinary first-order phase transition. In principle, the existence of critical points at the other crossover lines shown in Fig. 2 depends on the model parametrization. In this case, additional effects, such as non-equilibrium spinodal decomposition would have to be addressed. These effects have been explored, for example, in the context of Nambu–Jona-lasinio model [83, 84]. However, this is beyond the scope of the current work. We plan to elaborate on this interesting issue elsewhere.
@@ III Susceptibilities of the net-baryon number density
which contain contributions from the nucleonic, $\Delta$ sectors, and terms that mix them. We note that in [64], a simplified system consisting of a single parity doublet was considered. By explicitly including the $\Delta_{\pm}$ states, in addition to the individual nucleonic and $\Delta$ sectors, four additional susceptibilities that mix them have to be considered. The individual terms in Eq. (22) equation are given as follows
@@
The scaled variances, $R_{2,1}^{\alpha\beta}$, can be defined also for cumulants related to the nucleonic and $\Delta$ sectors in Eq. (24). We note that, in general, the partial scaled variances $R_{2,1}^{\alpha\beta}$, are not additive. For example, $R_{2,1}^{NN}+R_{2,1}^{\Delta\Delta}+2R_{2,1}^{N\Delta}\neq R_{2,1}^{B}$. We note that due to experimental limitations, the net-proton number fluctuations are assumed to reflect the fluctuations of the net-baryon number. However, this relation has not yet been explored in theoretical models with dynamical chiral symmetry restoration.
@@ IV Results
First, we study the susceptibilities at finite temperatures and vanishing baryon chemical potential. They are shown in Fig. 4, for $r_{\Delta}=0.8$. It is instructive to compare our results with those of the truncated hadron resonance gas (tHRG) model, where the only degrees of freedom are nucleons, $\Delta$s, and their chiral partners. The HRG model is widely used for the description of matter under extreme conditions, e.g., in the context of heavy-ion collision phenomenology [87, 88, 89, 90, 91, 92]. The thermodynamic potential of the tHRG model is a mixture of uncorrelated ideal gases of stable particles:
@@
We remark on the simplified nature the current model. Namely, the $N(1535)$, $\Delta(1232)$, and $\Delta(1700)$ resonances are considered to be stable particles by neglecting their finite widths. However, including the effects of finite width in a self-consistent way is non-trivial in the context of the relativistic mean-field approach. Forecasting possible effects on fluctuation observables would require to account for the imaginary part of the resonance self-energy as done, e.g., for $N_{-}(1535)$ in dense nuclear matter [93].
@@
Now we turn to the finite baryon chemical potential regime and consider fixed $T=0.03~{}\rm GeV$. In Fig. 5, we show the net-baryon number susceptibility. The behavior at $\mu_{B}<1~{}\rm GeV$ is similar for all the values of $r_{\Delta}$ shown in the figure. The curves show a well-pronounced peak which is a remnant of the liquid-gas phase transition. Small deviations from the pure nucleonic result are due to a small fraction of the $\Delta$ matter present at $T=0.03~{}$GeV. At higher chemical potentials, the susceptibilities develop plateaux. A small peak is seen for the pure nucleonic EOS around $\mu_{B}=1.4~{}\rm GeV$, which is a remnant of the first-order chiral phase transition at low temperatures. The secondary peak for $r_{\Delta}=0.8$ is a result of lower repulsion for the $\Delta$ matter, leading to an early onset of the $\Delta_{+}$ state. The result for $r_{\Delta}=1.1$ is very close to the pure nucleonic EOS. We note that for $r_{\Delta}=1.2$, the net-baryon number susceptibility converges to that of pure nucleonic EOS. This convergence is due to the suppression of $\Delta$ matter by repulsive interactions. However, the net-baryon number susceptibility at higher baryon chemical potential shows very little sensitivity to the choice of $r_{\Delta}$ and depends only on the composition of the asymptotic matter, as discussed in the previous section.
The individual susceptibilities $\chi_{2}^{\alpha\beta}$ also depend on the values of repulsive couplings. This can be seen in the right panel of Fig. 6. For $r_{\Delta}<1$, the susceptibility is dominated by the contribution of $\Delta_{\pm}$. Similarly, for $r_{\Delta}>1$, they are dominated by the contribution of $N_{\pm}$. This is expected from the relative abundances shown in Fig. 3 and discussed above. For $r_{\Delta}=1$ all susceptibilities are not negligible. In general, the susceptibilities $\chi_{2}^{\alpha\alpha}$ at high baryon chemical potential are much larger than the corresponding peak from the liquid-gas transition below $\mu_{B}=1~{}$GeV. At the same time, the positive-negative parity correlations within the same doublet become negative and show minima of similar size as the peaks in $\chi_{2}^{\alpha\alpha}$. The correlations between $N_{\pm}$ and $\Delta_{\pm}$ are shown in the left panel of Fig. 6 and have a qualitatively similar structure for different values of $r_{\Delta}$. However, their magnitude depends on the choice of the repulsive coupling and they are non-negligible, except for $r_{\Delta}=1.2$, for which the $\Delta$ matter is almost completely suppressed. Interestingly, we find that the susceptibilities of the opposite parity states (which are not chiral partners), i.e., $\chi_{2}^{N_{+}\Delta_{-}}$ and $\chi_{2}^{N_{-}\Delta_{+}}$ become large and negative. They become dominant and have minima around the chiral crossover, providing an additional source of cancellation of the net-baryon number fluctuations at finite density.
@@
As shown in Fig. 8, qualitatively similar structures can be seen in the scaled variances defined in Eqs. (27). At low chemical potential, the sensitivity to the liquid-gas transition is seen in $R_{2,1}^{N_{+}N_{+}}$. At high baryon chemical potential as the chiral symmetry becomes restored, the signal is dominated by the scaled variances of the negative parity states, which develop peaks related to the onset of the chiral partners, $N_{-}$ and $\Delta_{-}$. The mixed scaled variances between $N_{\pm}$ and $\Delta_{\pm}$ show patterns similar to the susceptibilities $\chi_{2}^{\alpha\beta}$. At high baryon chemical potential, the ratios $R_{2,1}^{N_{+}\Delta_{+}}$ and $R_{2,1}^{N_{-}\Delta_{-}}$ feature peaks related to the restoration of chiral symmetry, while $R_{2,1}^{N_{-}\Delta_{+}}$ and $R_{2,1}^{N_{+}\Delta_{-}}$ feature minima.
Most of the structure seen in the scaled variances is washed out when the full nucleonic and $\Delta$ sectors are considered as defined in Eqs. (22) and (27). This is shown in Fig. 9. For $r_{\Delta}=0.8$, $R_{2,1}^{NN}\simeq R_{2,1}^{B}$ below $\mu_{B}\simeq 0.9~{}\rm GeV$ until $\Delta$ matter becomes populated. This is also reflected in the enhancement of $R_{2,1}^{\Delta\Delta}$, which converges to $R_{2,1}^{B}$ with a higher baryon chemical potential. At the same time $R_{2,1}^{NN}$ goes to unity, due to $r_{\Delta}<1$. Similarly for $r_{\Delta}=1.0$, $R_{2,1}^{NN}$ deviates from $R_{2,1}^{B}$ as soon as the $\Delta$ matter sets in, i.e., $R_{2,1}^{\Delta\Delta}$ deviates from unity. However, since at high density the composition is determined by the degeneracy factors, none of the ratios converge to $R_{2,1}^{B}$. For $r_{\Delta}=1.2$, $R_{2,1}^{\Delta\Delta}\simeq 1$, due to the suppression of $\Delta_{\pm}$ states. Consequently, $R_{2,1}^{NN}\simeq R_{2,1}^{B}$. However, $R_{2,1}^{B}$ is suppressed compared to other scaled variances, regardless of the choice of $r_{\Delta}$.